# Probing Massive Fields with Multi-Band Gravitational-Wave Observations

### v1

@@ Abstract
We investigate the prospect of probing massive fields and testing gravitational theories with multi-band observations of gravitational waves emitted from coalescing compact binaries. Focusing on the dipole radiation induced by a massive field, we show that multi-band observations can probe the field with mass ranging from $10^{-16}{\rm eV}$ to $10^{-15}{\rm eV}$, a parameter space that cannot be probed by the milli-Hertz band observations alone. Multi-band observations can also improve the constraints obtained with the LIGO-Virgo-KAGRA binaries by at most 3 orders of magnitude in the mass range. Moreover, we show that multi-band observations can discriminate the spin of the field, which cannot be identified with single band observations.
@@ I Introduction
The existence of particles beyond standard model is not only predicted by fundamental theories, but also implied by observations of dark matter and dark energy. These new particles may weakly couple to the stand model particles, but could still get excited in extreme gravitational environments. It is known that, fields of the new particles, such as light axions, generalized Proca and extra degrees of freedom in gravity beyond general relativity (GR), can be significantly excited by neutron stars and black holes due to instability [1, 2, 3, 4, 5], nonlinearity [6, 7] and other mechanisms [8, 9, 10], providing promising ways of searching for physics beyond the standard model. For example, superradiantly excited axions around supermassive black holes can induce additional polarization of light, and hence are constrained with Event Horizon Telescope [11, 12].
@@
In recent years, gravitational waves (GWs) emitted by coalescing compact binaries have became a valuable probe to the physics in extreme gravitational environments. Deviations from GR and the standard model are searched and constrained by analyzing GW signals observed by the LIGO-Virgo-KAGRA (LVK) collaboration [13, 14, 15, 16, 17]. Also see, e.g., Refs. [18, 19, 20, 21, 22] and the references therein. With such a success in territorial detections, space-borne GW interferometers, such as LISA, Tianqin, and Taiji, are planned to launch by the mid-2030s. These space-borne detectors target GWs of milli-Hertz (mHz) frequencies, which are complementary to the LVK band, and hence will probe fundamental physics from different approaches.
@@
Interestingly, about $10$-$1000$ stellar mass black hole binaries are expected to be observed in both the mHz band and the LVK band [23], opening the prospect for multi-band GW astronomy. These binaries first inspiral in the mHz band for several years, and then re-appear in the LVK band typically a few weeks before they merge. Due to the long persistence of the signal in the mHz band, space-borne detectors can measure the masses and sky location of the binaries with great accuracy, while LVK and future territorial detectors can measure the GW amplitude better due to the high signal to noise ratio (SNR). Therefore, multi-band detection shall significantly improve parameter estimation in GW sources, and will be ideal for probing fundamental physics [24], such as constraining post-Newtonian (PN) and post-Einsteinian (PE) deviations [25, 26, 27, 28, 29], searching for dipole GW radiation [30, 31, 32, 33], measuring GW dispersion relation [34, 35], performing consistency tests [24, 36, 27] and bounding alternative gravity theories [36].
@@
In this work, we emphasize the prospect of probing new massive bosonic fields with multi-band GW observations. For massless (or ultralight) bosonic fields, their effects on orbital dynamics can be captured by parameterized PN and PE formalisms, and have been constrained with the Hulse-Taylor pulsar [37, 38, 39, 40], LVK binaries [17] and pulsar timing arrays[41]. Detectability of such fields with future GW detectors and multi-band observations is also forecasted in the literature [42, 43, 44, 45]. Effects of a massive field, however, do not generally fit with the parameterized PN or PE formalism, and should be treated separately. In particular, for fields with masses heavier than $10^{-16}{\rm eV}$, their effects are suppressed for inspirals in the mHz band [46, 47, 48, 44, 49]. Yet, they can be constrained with LVK binaries to certain accuracy, if their mass is below $10^{-11}{\rm eV}$ [50]. In this work, we shall demonstrate that constraints on massive fields can be significantly improved by 3 orders of magnitude with multi-band observations, especially for the fields with mass ranging in $10^{-16}{\rm eV}-10^{-15}{\rm eV}$, which cannot be well probed by space-born GW detectors. Moreover, we find that multi-band GW observations can distinguish the spin of the field, if its mass is within $10^{-16}-10^{-15}{\rm eV}$, which is hardly done with single band observations. We work in the units of $G=\hbar=c=1$.
@@ II Inspirals with massive fields
Though the excitation mechanism depends on the theories, the new field, once excited by the compact object, typically affects binary inspirals by mediating additional force and emitting additional radiation. While the force could modify the inspiral waveform at 00 PN order, the additional radiation usually manifests at $-1$ PN order, and hence is the main signature that we are after.
To be concrete, we start from a massive spin-0 field with following action,
@@
where $\varphi$ is the spin-0 field of mass $\mu$, and $\mathcal{L}_{\rm int}$ denotes the interaction between the field and the matter fields. During early inspiral, the two compact objects in a binary can be treated as two non-relativistic point-like particles. In this case, the leading order interaction term is
@@
where $Q_{i}$ and ${\bf x}_{i}(t)$ represent the charge and position of the $i$-th object respectively. We shall also consider a massive spin-1 field, the action of which is
@@
where $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$. The spin-1 field couples to matter fields through their current $J^{\mu}$, i.e., $\mathcal{L}_{\rm int}\propto A_{\mu}J^{\mu}$. Again, treating the binary as two non-relativistic point-like particles, the coupling becomes
@@
In principle, one can consider massive spin-2 field which presents, for example, in bi-gravity. However, defining energy flux is subtle in theories with two dynamical metrics [51]. It is also possible that the graviton itself has a non-zero but tiny mass, and GR should be replaced by a massive gravity theory. In this case, nonlinearity is expected to be important within the so-call Vainshtein radius, a scale that is typically much larger than the size of stellar mass binary. As a results, deviations from GR, including radiations of extra degrees of freedom in massive gravity, are expected to be suppressed for stellar mass binaries. For these reasons, we shall focus on and demonstrate the multi-band detection strategy with only spin-0 and spin-1 fields.
@@
where $R$ is the orbital separation, $\Omega$ is the orbital frequency and $\Delta q\equiv\left(Q_{1}/m_{1}\right)-\left(Q_{2}/m_{2}\right)$ with $m_{1,2}$ being the mass of compact objects. See, e.g., Ref. [52] for details. The models we considered here is generic, and hence the energy flux calculated by Eq. (5) applies to many theories, see Refs. [53, 54, 55, 56, 57, 44, 58, 59, 60] for example.
To incorporate the effects of the dipole radiation from the massive fields into inspiral waveform, we calculate the waveform in frequency domain by extending the TaylorF2 template. Specifically, the waveform template is given in frequency domain,
@@
where $\Psi_{\rm GR}$ is the GW phase predicted by GR, and the explicit expression of $\Psi_{\rm MF}$ is given in App. A.
Figure 1: Detectability on spin-0 and spin-1 fields of different mass. The markers show the expectation errors on $\Delta q$ obtained by calculating the Fisher information matrix (cf. Eq. (19)), assuming a GW150914-like event observed in the mHz and the LVK bands. Here circles and triangles show the errors for spin-0 and spin-1 fields respectively, while blue and red denote the mHz and the LVK bands. For observations in the mHz bands, markers jointed by the solid, dashed and dotted lines show the situation of a $4$-year observation starting at $4$ years before merger, a $1$-year observation starting at $1$ year before merger and a $1$-year observation starting at $4$ years before merger respectively. The green lines show the merger time deviation from GR, when there is a spin-0 or a spin-1 massive field.
@@ III Single-band observations
Given the inspiral waveform, we first investigate detectability of the massive fields with single band observations. It is convenient to introduce
@@
with $M=m_{1}+m_{2}$ and $\nu=m_{1}m_{2}/M^{2}$. Here $f$ should be $f_{\mu}$ or $f_{i}$, whichever is larger, with $f_{i}\sim 2{\rm Hz}$ being the GW frequency when the inspiral signal enters the LVK band. Assuming the dipole radiation can be detected if $\Psi_{\rm MF}\sim{\cal O}(1)$, we expect that detectors in the LVK band are sensitive to $\Delta q\gtrsim 0.008$ for fields with $f_{\mu}<f_{i}$, and the sensitivity gets worse as $f_{\mu}$ increases. We shall also estimate the detectability with observations in the mHz band. Different from the LVK band, we may not observe a clear chirping in GW frequncy during the observation time, if the binary is in its very early inspiral stage when the detector turns on. In particular, the time to merger given by GR is
@@
Therefore, if the binary is emitting GWs with $f_{i}\gtrsim 0.03{\rm Hz}$ when detector turns on, the signal will leave the mHz band during the observation period, which is typically assumed to be $1$ year or $4$ years. The phase shift induced by the dipole radiation during the observation time can still be estimated with Eq. (12), and we expect the minimally detectable $\Delta q$ is $\sim 6\times 10^{-5}$ for $f_{\mu}<f_{i}$. On the other hands, for $f_{i}\ll 0.03{\rm Hz}$, we do not expect to see $f$ changes significantly during the operating time, and the detector can only probe the dipole radiation of fields with $f_{\mu}\lesssim f_{i}<0.03{\rm Hz}$. In this case, the additional phase shift is given by
@@
where $T_{\rm obs}$ is operating time. The minimally detectable $|\Delta q|$ is $\sim 3\times 10^{-5}$ assuming a $4$-year observation. Comparing to the LVK band, observations in the mHz band generally have better accuracy for fields with $f_{\mu}\lesssim 0.01{\rm Hz}$. However, the accuracy gets worse quickly as $f_{\mu}$ increases. In fact, observations in the mHz band almost cannot probe fields with $f_{\mu}>0.1{\rm Hz}$, because the massive field almost plays no role during the entire observation.
A rigorous forecast on detectability can be made using Fisher information matrix. Having in mind an interferometric detector and working in frequency domain, the measured date $d$ can be expressed as a linear combination of the signal $s$ and the detector noise $n$,
@@
For demonstration, we consider a GW150914-like event, i.e., a binary with $m_{1}=36M_{\odot}$ and $m_{2}=29M_{\odot}$ inspiralling at $410{\rm Mpc}$, and estimate the detectability of $|\Delta q|$ by calculating the Fisher information matrix for a spin-0 field with certain $\mu$. For simplicity, we average over the binary’s sky-location, inclination and polarization, and ignore the spin effects for simplicity. Then the parameters reduce to
@@
To forecast the detectability in mHz band, we use the effective sensitivity curve of LISA [62] obtained after averaging over sky and polarization angle. Given an GW150914-like event, we may have different situations, and the detectability on $\Delta q$ is shown in Fig. 1. We first consider a $1$-year observation, starting at $1$ year before the binary merges. In this case, we have $f_{i}\simeq 0.029{\rm Hz}$. With such an observation, LISA can detect $\Delta q>2\times 10^{-6}$ for $f_{\mu}<0.029{\rm Hz}$. We then consider a $4$-year observation, starting at $4$ years before the binary merges, in which case $f_{i}\simeq 0.017{\rm Hz}$ and LISA can detect $\Delta q>8\times 10^{-7}$ for $f_{\mu}<0.017{\rm Hz}$. We find that, for $f_{\mu}\ll f_{i}$, the detectability approximately improves with the observation duration, which is expected from Eq. (14). In both cases, the detectability becomes worse quickly as $f_{\mu}$ approaches $0.1{\rm Hz}$, as beyond which the field barely affects the orbital dynamics in the observation period. In particular, detectability of $1$-year and $4$-year observations is almost the same for $f_{\mu}>0.02{\rm Hz}$, because the massive field only become dynamically relevant in the last year of the observation. For comparison, we also consider a $1$-year observation, but starting at $4$ years before the binary merges. Given such an observation, LISA can detect $\Delta q>2\times 10^{-6}$ for $f_{\mu}<0.01{\rm Hz}$, and quickly loses the detectability as $f_{\mu}$ approaches to $0.017{\rm Hz}$, as beyond which the mass field becomes irrelevant to inspiral during the observation period. We perform a similar analysis for spin-1 field as well, and show the results in Fig. 1.
IV Multi-band observations
@@
Figure 2: Parameter constraints for a GW150914-like event, i.e., $m_{1}=36M_{\odot}$, $m_{2}=29M_{\odot}$ and $d_{L}=410{\rm Mpc}$, obtained with mHz band (blue) and LVK band (red) observations. When calculating the Fisher information matrix, we assume the presence of a spin-0 field with $f_{\mu}=1{\rm Hz}$.
@@
Figure 3: Discrimination on the spin of massive fields. The blue (red) lines show the difference between the spin-0 and the spin-1 field in GW phase, $\Delta\Psi\equiv\left|\Psi_{\rm MF}^{s=0}-\Psi_{\rm MF}^{s=1}\right|$, assuming a GW150914-like event observed in mHz (LVK) band. The green lines show the difference between the spin-0 and the spin-1 field in merger time $\Delta t$. The green region shows the parameter space where the detectability on the spin of the field can be improved by multi-band observations.
@@
As demonstrated previously, observations in the mHz band alone can improve the detectability on massive fields, but only if for fields with $f_{\mu}<0.1{\rm Hz}$. Nevertheless, fields with $f_{\mu}>0.1{\rm Hz}$ can still affect the later binary evolution and accelerate orbital decay while observations in the mHz band can forecast the merger time with an accuracy of ${\cal O}(1)$ second. With a multi-band observation such that the binary is also observed later in the LVK band, the merger time forecasted by the mHz observations can be tested in the LVK band, providing an additional constraint on the dipole radiation induced by the massive fields. In Fig. 2, we consider a GW150914-like event, and estimate the constraints of the waveform parameters assuming there is a spin-0 field with $f_{\mu}=1{\rm Hz}$. Given the mass of the field, LVK band observation can probe the field if $|\Delta q|>10^{-2}$, while the mHz band observation alone cannot probe the filed at all. For the other parameters, observation in LVK band does better measurement on distance and merger time, while observation in mHz band does better measurement on masses. In particular, the mHz band can measure the merger time with an accuracy of $0.03s$ at $90\%$ confidence level.
@@
Assuming $|\Delta q|\ll 1$, the merger time in the presence of dipole radiation will be earlier than that in GR by amount of
@@
for $0.1{\rm Hz}<f_{\mu}<2{\rm Hz}$. In Fig. 1, we show the constraints on $|\Delta q|$ for fields with different masses by requiring $\Delta t<0.03{\rm s}$, assuming a multi-band observation of GW150914-like event. Comparing to the single band observations, we find that the multi-band observation indeed improves the detectability on $\Delta q$ by filling the gap between the mHz and the LVK bands.
Beside improving detectability of the massive fields, multi-band observation can further distinguish spin-0 fields from spin-1 fields. With observations in the LVK band or in the mHz band alone, we expect to distinguish the fields if they could induce ${\cal O}(1)$ difference in GW phase, namely $\left|\Psi_{\rm MF}^{s=0}-\Psi_{\rm MF}^{s=1}\right|\sim{\cal O}(1)$. Given Eq. (5), waveforms of spin-0 and spin-1 field differ notably only when the dipole radiation just turns on. Therefore, if $\mu\ll\Omega$ during the single band observation, one cannot distinguish whether the dipole radiation is caused by a spin-0 or spin-1 field, cf. Eq. (5), even with a nontrivial measurement of $\Delta q$. With multi-band observations, one can further infer whether the dipole radiation observed in the LVK band is from a spin-0 or spin-1 field by examine the merger time. Considering the GW150914-like event, Fig. 3 shows the GW phase difference for observations in the LVK and the mHz band, as well as the difference in merger time $\Delta t=|t_{c}^{s=0}-t_{c}^{s=1}|$, from which we can conclude that multi-band observations can distinguish the spin of the fields in some of the parameter space where the single-band observation cannot.

### v3

# Probing massive fields with multiband gravitational-wave observations

@@ Abstract
We investigate the prospect of probing massive fields and testing gravitational theories with multiband observations of gravitational waves emitted from coalescing compact binaries. Focusing on the dipole radiation induced by a massive field, we show that multiband observations can probe the field with mass ranging from $10^{-16}$ to $10^{-15}{\rm eV}$, a parameter space that cannot be probed by the millihertz band observations alone. Multiband observations can also improve the constraints obtained with the LIGO-Virgo-KAGRA binaries by at most 3 orders of magnitude in the mass range. Moreover, we show that multiband observations can discriminate the spin of the field, which cannot be identified with single-band observations.
@@ I Introduction
The existence of particles beyond the standard model is not only predicted by fundamental theories, but also implied by observations of dark matter and dark energy. These new particles may weakly couple to the standard model particles, but could still get excited in extreme gravitational environments. It is known that fields of the new particles, such as light axions, generalized Proca, and extra degrees of freedom in gravity beyond general relativity (GR), can be significantly excited by neutron stars and black holes due to instability [1, 2, 3, 4, 5], nonlinearity [6, 7], and other mechanisms [8, 9, 10], providing promising ways of searching for physics beyond the standard model. For example, superradiantly excited axions around supermassive black holes can induce additional polarization of light, and hence are constrained with the Event Horizon Telescope [11, 12].
@@
In recent years, gravitational waves (GWs) emitted by coalescing compact binaries have became a valuable probe to the physics in extreme gravitational environments. Deviations from GR and the standard model are searched and constrained by analyzing GW signals observed by the LIGO-Virgo-KAGRA (LVK) Collaboration [13, 14, 15, 16, 17] (also see, e.g., Refs. [18, 19, 20, 21, 22] and the references therein). With such a success in territorial detections, space-borne GW interferometers, such as LISA, TianQin, and Taiji, are planned to launch by the mid-2030s. These space-borne detectors target GWs of millihertz frequencies, which are complementary to the LVK band, and hence will probe fundamental physics from different approaches.
@@
Interestingly, about 101010101000100010001000 stellar mass black hole binaries are expected to be observed in both the millihertz band and the LVK band [23], opening the prospect for multiband GW astronomy. These binaries first inspiral in the millihertz band for several years, and then reappear in the LVK band typically a few weeks before they merge. Because of the long persistence of the signal in the millihertz band, space-borne detectors can measure the masses and sky location of the binaries with great accuracy, while LVK and future territorial detectors can measure the GW amplitude better due to the high signal-to-noise ratio (SNR). Therefore, multiband detection shall significantly improve parameter estimation in GW sources, and will be ideal for probing fundamental physics [24], such as constraining post-Newtonian (PN) and post-Einsteinian (PE) deviations [25, 26, 27, 28, 29], searching for dipole GW radiation [30, 31, 32, 33], measuring GW dispersion relation [34, 35], performing consistency tests [24, 36, 27] and bounding alternative gravity theories [36].
In this work, we emphasize the prospect of probing new massive bosonic fields with multiband GW observations. For massless (or ultralight) bosonic fields, their effects on orbital dynamics can be captured by parametrized PN and PE formalisms and have been constrained with the Hulse-Taylor pulsar [37, 38, 39, 40], LVK binaries [17] and pulsar timing arrays [41]. They can also be probed by observations of GWs from extreme mass-ratio inspirals in the millihertz band [42].
@@
Detectability of such fields with future GW detectors and multiband observations is also forecasted in the literature [43, 44, 45, 46]. Effects of a massive field, however, do not generally fit with the parametrized PN or PE formalism and should be treated separately. In particular, for fields with masses heavier than $10^{-16}\,{\rm eV}$, their effects are suppressed for inspirals in the millihertz band [47, 48, 49, 45, 50, 51]. Nevertheless, they can be constrained with LVK binaries to certain accuracy, if their mass is below $10^{-11}\,{\rm eV}$ [52]. In this work, we shall demonstrate that constraints on massive fields can be significantly improved by 3 orders of magnitude with multiband observations, especially for the fields with mass ranging in $10^{-16}\text{--}10^{-15}\,{\rm eV}$, which cannot be well probed by space-borne GW detectors. Moreover, we find that multiband GW observations can distinguish the spin of the field, if its mass is within $10^{-16}\text{--}10^{-15}\,{\rm eV}$, which is hardly done with single-band observations. We work in the units of $G=\hbar=c=1$.
@@ II Inspirals with massive fields
Though the excitation mechanism depends on the theories, the new field, once excited by the compact object, typically affects binary inspirals by mediating additional force and emitting additional radiation. While the force could modify the inspiral waveform at 00 PN order, the additional radiation usually manifests at $-1$ PN order [48], and hence is the main signature that we are after.
To be concrete, we start from a massive spin-0 field with following action:
@@
where $\varphi$ is the spin-0 field of mass $\mu$, and $\mathcal{L}_{\rm int}$ denotes the interaction between the field and the matter fields. During early inspiral, the two compact objects in a binary can be treated as two nonrelativistic pointlike particles. In this case, the leading order interaction term is
@@
where $Q_{i}$ and ${\bf x}_{i}(t)$ represent the charge and position of the $i$th object respectively. We shall also consider a massive spin-1 field, the action of which is
@@
where $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$. The spin-1 field couples to matter fields through their current $J^{\mu}$, i.e., $\mathcal{L}_{\rm int}\propto A_{\mu}J^{\mu}$. Again, treating the binary as two nonrelativistic pointlike particles, the coupling becomes
@@
In principle, one can consider massive spin-2 field which presents, for example, in bigravity. However, defining energy flux is subtle in theories with two dynamical metrics [53]. It is also possible that the graviton itself has a nonzero but tiny mass, and GR should be replaced by a massive gravity theory. In this case, nonlinearity is expected to be important within the so-called Vainshtein radius, a scale that is typically much larger than the size of stellar mass binary. As a results, deviations from GR, including radiations of extra degrees of freedom in massive gravity, are expected to be suppressed for stellar mass binaries. For these reasons, we shall focus on and demonstrate the multiband detection strategy with only spin-0 and spin-1 fields.
@@
where $R$ is the orbital separation, $\Omega$ is the orbital frequency and $\Delta q\equiv\left(Q_{1}/m_{1}\right)-\left(Q_{2}/m_{2}\right)$ with $m_{1,2}$ being the mass of compact objects (see, e.g., Ref. [54] for details). The models we consider here are generic, and hence the energy flux calculated by Eq. (5) applies to many theories: see Refs. [55, 56, 57, 58, 59, 45, 60, 61, 62, 63, 64] for example.
To incorporate the effects of the dipole radiation from the massive fields into inspiral waveform, we calculate the waveform in frequency domain by extending the TaylorF2 template. Specifically, the waveform template is given in the frequency domain,
@@
where $\Psi_{\rm GR}$ is the GW phase predicted by GR, and the explicit expression of $\Psi_{\rm MF}$ is given in the Appendix. A.
Figure 1: Detectability on spin-0 and spin-1 fields of different mass. The markers show the expectation errors on $\Delta q$ obtained by calculating the Fisher information matrix [cf. Eq. (19)], assuming a GW150914-like event observed in the millihertz and the LVK bands. Here circles and triangles show the errors for spin-0 and spin-1 fields, respectively, while blue and red denote the millihertz and the LVK bands. For observations in the millihertz bands, markers jointed by the solid, dashed, and dotted lines show the situation of a $4$-yr observation starting at four years before merger, a $1$-yr observation starting at one year before merger, and a $1$-yr observation starting at four years before merger, respectively. The green lines show the merger time deviation from GR, when there is a spin-0 or a spin-1 massive field.
@@ III Single-band observations
Given the inspiral waveform, we first investigate detectability of the massive fields with single-band observations. It is convenient to introduce
@@
with $M=m_{1}+m_{2}$ and $\nu=m_{1}m_{2}/M^{2}$. Here $f$ should be $f_{\mu}$ or $f_{i}$, whichever is larger, with $f_{i}\sim 2\,{\rm Hz}$ being the GW frequency when the inspiral signal enters the LVK band. Assuming the dipole radiation can be detected if $\Psi_{\rm MF}\sim{\cal O}(1)$, we expect that detectors in the LVK band are sensitive to $\Delta q\gtrsim 0.008$ for fields with $f_{\mu}<f_{i}$, and the sensitivity gets worse as $f_{\mu}$ increases. We shall also estimate the detectability with observations in the millihertz band. Different from the LVK band, we may not observe a clear chirping in GW frequency during the observation time, if the binary is in its very early inspiral stage when the detector turns on. In particular, the time to merger given by GR is
@@
Therefore, if the binary is emitting GWs with $f_{i}\gtrsim 0.03\,{\rm Hz}$ when detector turns on, the signal will leave the millihertz band during the observation period, which is typically assumed to be $1$ or $4$ yr. The phase shift induced by the dipole radiation during the observation time can also be estimated with Eq. (12), and we expect the minimally detectable $\Delta q$ is $\sim 6\times 10^{-5}$ for $f_{\mu}<f_{i}$. On the other hand, for $f_{i}\ll 0.03\,{\rm Hz}$, we do not expect to see $f$ changing significantly during the operating time, and the detector can only probe the dipole radiation of fields with $f_{\mu}\lesssim f_{i}<0.03\,{\rm Hz}$. In this case, the additional phase shift is given by
@@
where $T_{\rm obs}$ is operating time. The minimally detectable $|\Delta q|$ is $\sim 3\times 10^{-5}$ assuming a $4$-yr observation. Comparing to the LVK band, observations in the millihertz band generally have better accuracy for fields with $f_{\mu}\lesssim 0.01\,{\rm Hz}$. However, the accuracy gets worse quickly as $f_{\mu}$ increases. In fact, observations in the millihertz band almost cannot probe fields with $f_{\mu}>0.1\,{\rm Hz}$, because the massive field almost plays no role during the entire observation.
A rigorous forecast on detectability can be made using Fisher information matrix. Having in mind an interferometric detector and working in frequency domain, the measured data $d$ can be expressed as a linear combination of the signal $s$ and the detector noise $n$,
@@
For demonstration, we consider a GW150914-like event, i.e., a binary with $m_{1}=36M_{\odot}$ and $m_{2}=29M_{\odot}$ inspiralling at $410\,{\rm Mpc}$, and estimate the detectability of $|\Delta q|$ by calculating the Fisher information matrix for a spin-0 field with certain $\mu$. For simplicity, we average over the binary’s sky location, inclination, and polarization, and ignore the spin effects for simplicity. Then the parameters reduce to
@@
To forecast the detectability in the millihertz band, we use the effective sensitivity curve of LISA [66] obtained after averaging over sky and polarization angles. Given an GW150914-like event, we may have different situations, and the detectability on $\Delta q$ is shown in Fig. 1. We first consider a $1$-yr observation, starting at $1$ yr before the binary merges. In this case, we have $f_{i}\simeq 0.029\,{\rm Hz}$. With such an observation, LISA can detect $\Delta q>2\times 10^{-6}$ for $f_{\mu}<0.029\,{\rm Hz}$. We then consider a $4$-yr observation, starting at $4$ yr before the binary merges, in which case $f_{i}\simeq 0.017\,{\rm Hz}$ and LISA can detect $\Delta q>8\times 10^{-7}$ for $f_{\mu}<0.017\,{\rm Hz}$. We find that, for $f_{\mu}\ll f_{i}$, the detectability approximately improves with the observation duration, which is expected from Eq. (14). In both cases, the detectability becomes worse quickly as $f_{\mu}$ approaches $0.1\,{\rm Hz}$, as beyond which the field barely affects the orbital dynamics in the observation period. In particular, detectability of $1$-and $4$-yr observations is almost the same for $f_{\mu}>0.02\,{\rm Hz}$, because the massive field only becomes dynamically relevant in the last year of the observation. For comparison, we also consider a $1$-yr observation, but starting at $4$ yr before the binary merges. Given such an observation, LISA can detect $\Delta q>2\times 10^{-6}$ for $f_{\mu}<0.01\,{\rm Hz}$, and quickly loses the detectability as $f_{\mu}$ approaches to $0.017\,{\rm Hz}$, as beyond which the mass field becomes irrelevant to inspiral during the observation period. We perform a similar analysis for a spin-1 field as well and show the results in Fig. 1.
IV multiband observations
@@
Figure 2: Parameter constraints for a GW150914-like event, i.e., $m_{1}=36M_{\odot}$, $m_{2}=29M_{\odot}$ and $d_{L}=410{\rm Mpc}$, obtained with millihertz band (blue) and LVK band (red) observations. When calculating the Fisher information matrix, we assume the presence of a spin-0 field with $f_{\mu}=1\,{\rm Hz}$.
@@
Figure 3: Discrimination on the spin of massive fields. The blue (red) lines show the difference between the spin-0 and the spin-1 fields in GW phase, $\Delta\Psi\equiv\left|\Psi_{\rm MF}^{s=0}-\Psi_{\rm MF}^{s=1}\right|$, assuming a GW150914-like event observed in the millihertz (LVK) band. The green line shows the difference between the spin-0 and the spin-1 field in merger time $\Delta t$. Region above the green line and below the blue and red lines is the parameter space where the detectability on the spin of the field is expected to be improved by multiband observations.
@@
As demonstrated previously, observations in the millihertz band alone can improve the detectability on massive fields, but only for fields with $f_{\mu}<0.1\,{\rm Hz}$. Nevertheless, fields with $f_{\mu}>0.1\,{\rm Hz}$ can still affect the later binary evolution and accelerate orbital decay while observations in the millihertz band can forecast the merger time with an accuracy of ${\cal O}(1)$ s. With a multiband observation that the binary is observed later in the LVK band, the merger time forecasted by the millihertz observations can be tested in the LVK band, providing an additional constraint on the dipole radiation induced by the massive fields. In Fig. 2, we consider a GW150914-like event and estimate the constraints of the waveform parameters assuming there is a spin-0 field with $f_{\mu}=1\,{\rm Hz}$. The constraints are obtained using the Fisher information matrix as stated in Sec. III. Given the mass of the field, LVK band observation can probe the field if $|\Delta q|>10^{-2}$, while the millihertz band observation alone cannot probe the field at all. For other parameters, observation in the LVK band does better measurement on distance and merger time, while observation in the millihertz band does better measurement on masses. In particular, the millihertz band can measure the merger time with an accuracy of $0.03$  s at $90\%$ confidence level.
@@
Assuming $|\Delta q|\ll 1$, the merger time in the presence of dipole radiation will be earlier than that in GR by an amount of
@@
for $0.1<f_{\mu}<2\,{\rm Hz}$. In Fig. 1, we show the constraints on $|\Delta q|$ for fields with different masses by requiring $\Delta t<0.03\,{\rm s}$, assuming a multiband observation of a GW150914-like event. Comparing to the single-band observations, we find that the multiband observation indeed improves the detectability on $\Delta q$ by filling the gap between the millihertz and the LVK bands.
In addition to improving detectability of the massive fields, multiband observation can further distinguish spin-0 fields from spin-1 fields. With observations in the LVK band or in the millihertz band alone, we expect to distinguish the fields if they could induce ${\cal O}(1)$ difference in GW phase, namely, $\left|\Psi_{\rm MF}^{s=0}-\Psi_{\rm MF}^{s=1}\right|\sim{\cal O}(1)$. Given Eq. (5), waveforms of spin-0 and spin-1 fields differ notably only when the dipole radiation just turns on. Therefore, if $\mu\ll\Omega$ during the single-band observation, one cannot distinguish whether the dipole radiation is caused by a spin-0 or spin-1 field, cf. Eq. (5), even with a nontrivial measurement of $\Delta q$. With multiband observations, one can further infer whether the dipole radiation observed in the LVK band is from a spin-0 or spin-1 field by examining the merger time. Considering the GW150914-like event, Fig. 3 shows the GW phase difference for observations in the LVK and the millihertz band, as well as the difference in merger time $\Delta t=|t_{c}^{s=0}-t_{c}^{s=1}|$, from which we can conclude that multiband observations can distinguish the spin of the fields in some of the parameter space where the single-band observation cannot.